40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 8 08:08:18 2019, gautam, Update, PEM, PEM BLRMS anomaly PEManomaly.png
    Reply  Wed Oct 23 11:49:21 2019, gautam, Update, PEM, PEM BLRMS anomaly seisAll_20191021.pdfspecGrams.zip
       Reply  Thu Oct 24 18:37:15 2019, gautam, Update, PEM, T240 checkout 
          Reply  Tue Nov 5 12:37:50 2019, gautam, Update, PEM, T240 interface unit pulled out T240_recenter.pdfffPotential.pdf
             Reply  Wed Nov 13 19:34:45 2019, gautam, Update, PEM, Follow-up on seismometer discussion seisAll_20191111.pdfffPotential.pdf
                Reply  Wed Nov 13 20:15:56 2019, rana, Update, PEM, Follow-up on seismometer discussion 
                   Reply  Wed Nov 13 23:40:15 2019, gautam, Update, PEM, Follow-up on seismometer discussion seisAll_20191111_1Hz.pdf
                      Reply  Thu Nov 14 12:11:04 2019, rana, Update, PEM, Follow-up on seismometer discussion 
                      Reply  Fri Nov 15 00:18:41 2019, rana, Update, PEM, Follow-up on seismometer discussion MCshake.pdf
                         Reply  Fri Nov 15 12:16:48 2019, gautam, Update, PEM, Follow-up on seismometer discussion BS_ZspecGram.pdf.zip
                            Reply  Mon Nov 18 14:32:53 2019, gautam, Update, PEM, Follow-up on seismometer discussion 
                               Reply  Tue Nov 19 21:53:57 2019, gautam, Update, PEM, Follow-up on seismometer discussion seisxSpec.pdf
Message ID: 15027     Entry time: Fri Nov 15 00:18:41 2019     In reply to: 15024     Reply to this: 15030
Author: rana 
Type: Update 
Category: PEM 
Subject: Follow-up on seismometer discussion 

The large ground motion at 1 Hz started up again tonight at around 23:30. I walked around the lab and nearby buildings with a flashlight and couldn't find anything whumping. The noise is very sinusoidal and seems like it must be a 1 Hz motor rather than any natural disturbance or traffic, etc. Suspect that it is a pump in the nearby CES building which is waking up and running to fill up some liquid level. Will check out in the morning.

Estimate of displacement noise based on the observed MC_F channel showing a 25 MHz peak-peak excursion for the laser:

dL = 25e6 * (13 m / (c / lambda)

      = 1 micron

So this is a lot. Probably our pendulum is amplifying the ground motion by 10x, so I suspect a ground noise of ~0.1 micron peak-peak.

(this is a native PDF export using qtgrace rather than XMgrace. uninstall xmgrace and symlink to qtgrace.)

Attachment 1: MCshake.pdf  1.803 MB  | Hide | Hide all
MCshake.pdf
ELOG V3.1.3-